OpenStudy (anonymous):

let f be twice differentiable with f(0)=6, f(1)=5, and f'(1)=2. Evaluate the integral (x(f^n)x dx) from 0 to 1

OpenStudy (anonymous):

\[\int\limits xf^{n}(x)dx\]???

OpenStudy (amistre64):

you forgot the bounds :)

OpenStudy (amistre64):

\[\int^{1}_{0}x\ f''(x)\ dx\]

OpenStudy (amistre64):

\(x\ f'(x) -f(x)\) if we do ibps

OpenStudy (amistre64):

\[x\ f'(x) -f(x)-(x\ f'(x) -f(x))\] \[1\ f'(1) -f(1)-(0\ f'(0) -f(0))\] \[1\ (2) -5-(0\ f'(0) -6)\] \[2-5+6=3\]

OpenStudy (amistre64):

thats my story and im sticking to it :)

OpenStudy (anonymous):

is actually f to the power of n

OpenStudy (anonymous):

how do you get the integral distribute

OpenStudy (amistre64):